OpenStudy (rayep12):

Write the equation of a line that is perpendicular to y=-1 and that passes through the point (8, -4)

OpenStudy (acespeedfighter):

DO IT YOURSELF YOU SENT THIS MSG LIKE 100 TIMES

OpenStudy (rayep12):

i have one more after this

OpenStudy (welshfella):

slope of the required line = -1/-1

OpenStudy (welshfella):

call this slope m then use y - y1 = m(x-x1) to find its equation

OpenStudy (triciaal):

@welshfella slope of the required line = -1/0 so m = 0 @rayep12 use the formula given by @welshfella your (x1, y1) is the point (8, -4)

OpenStudy (welshfella):

OH Yes Sorry that was a mistake. The slope of the line y = -1 is not -1!!!

OpenStudy (welshfella):

in fact the slope of the line we need to find is infinite so its a vertical line

OpenStudy (welshfella):

passing through (8,-4)

OpenStudy (welshfella):

you dont need to use the formula I gave

OpenStudy (triciaal):

@rayep12 do you know how to find the slope of a line? can you find the equation of the line using the formula given? do you need more help? any questions? don't know where you are and you provide some input?

OpenStudy (triciaal):

correction can you provide some input?